### v1

# Classification of same-gate quantum circuits and their space-time symmetries with application to the level-spacing distribution

Abstract
We study Floquet systems with translationally invariant nearest-neighbor 2-site gates. Depending on the order in which the gates are applied on an $N$-site system with periodic boundary conditions, there are factorially many different circuit configurations. We prove that there are only $N-1$ different spectrally equivalent classes which can be viewed either as a generalization of the brick-wall or of the staircase configuration. Every class, characterized by two integers, has a nontrivial space-time symmetry with important implications for the level-spacing distribution – a standard indicator of quantum chaos. Namely, in order to study chaoticity one should not look at eigenphases of the Floquet propagator itself, but rather at the spectrum of an appropriate root of the propagator.
Introduction.– Chaoticity and integrability are important theoretical notions. Integrability can allow for analytical results, while chaotic systems, in spite of unpredictability of trajectories, adhere to statistical laws. So-called toy models – the simplest models with a given property – play an important role. Classical single-particle $H$ in one dimension (1D) is always integrable; one needs at least a 3D phase space for chaos to be possible. This can also be achieved already in 1D Ott (1993) by a time-dependent $H(t)$, the simplest case being a “kicked” system of form $H(t)=\frac{p^{2}}{2}+V(q)\tau\delta_{\tau}(t)$, where $\delta_{\tau}(t)$ is a train of delta functions. A canonical example is the standard map Casati et al. (1979). Similar logic of taking a Floquet propagator (map) works for quantum systems as well. For instance, one can exemplify single-particle quantum chaos with a kicked top Haake (2010); Haake et al. (1987). Going to many-body quantum systems a plethora of possibilities opens up, one choice being a Floquet propagator that is a product of two simpler propagators, e.g. Ref. Prosen (1999). In light of experimental advances in noisy quantum simulations Bloch et al. (2008); Mi et al. (2021); Kim et al. (2023) it pays off to consider systems where the basic building block is a nearest-neighbor gate rather than the local Hamiltonian (applying two-site gates is also simpler in classical numerical simulations Schollwöck (2011)).
We therefore focus on circuits where a one-unit-of-time Floquet propagator $F$ is composed of applying the same two-site unitary gate $V$ on all nearest-neighbor pairs of qubits in 1D – we call such systems simple circuits. Simple circuits have translational and temporal invariance, and, depending on the chosen gate $V$, span all dynamical regimes from integrability to chaos. Needles to say, such simple circuits have been extensively studied, a non-exhaustive list of only few of recent works includes Refs. Akila et al. (2016); Gritsev and Polkovnikov (2017); Ljubotina et al. (2019); Gopalakrishnan and Lamacraft (2019); Zhou and Nahum (2020); Klobas et al. (2021); Bertini et al. (2019); Chan et al. (2022); Nahum et al. (2022); Claeys et al. (2022); Žnidarič (2023). We show that any such circuit has a very simple form: it is a product of a local 2-site transformation and a free propagation (translation), and that there are only $N-1$ spectrally inequivalent classes.
There are many indicators of quantum chaos with perhaps the most frequently used one being the so-called level spacing distribution (LSD) $P(s)$ of nearest-neighbor eigenenergy spacing $s$ Haake (2010). According to the quantum chaos conjecture Bohigas et al. (1984) Hamiltonian systems with a chaotic classical limit  [^1] For many-body quantum systems a word of caution is in place. Because the level spacing is exponentially small in $N$ the LSD probes “chaoticity” on an exponentially large timescale. There can be situations where the relevant physics is not chaotic at all despite the RMT LSD, see e.g. Ref. Santos (2004); Brenes et al. (2018) are expected to display $P(s)$ given by the random matrix theory (RMT) Mehta (2004). For Floquet systems checking for quantum chaos via $P(s)$ is even simpler: writing eigenvalues of $F$ as ${\rm e}^{{\rm i}\phi_{j}}$ the density of eigenphases $\phi_{j}$ should be uniform and therefore taking for $s=(\phi_{j+1}-\phi_{j}){\cal N}/2\pi$ there is no need for the unfolding that is required when studying spectra of $H$ Haake (2010). It is therefore rather surprising that, while there are hundreds of papers using $P(s)$ to study chaoticity in many-body Hamiltonian systems, there are essentially none studying LSD in simple (same-gate) quantum circuits (exceptions are recent Refs. Duh ; Huang et al. (2023)). The reason is that, surprisingly, LSD for chaotic simple circuits seemingly does not adhere to the RMT expectation. In our paper we will show that the reason behind it is a space-time symmetry that all such circuits posses.
Let us demostrate that by a simple generalized brick-wall (BW) circuit with 3 layers (Fig. 1 inset), where we translate each 2-site gate by 3 sites (instead of 2 as in BW). For periodic boundary conditions and $N$ divisible by $3$ the Floquet propagator can be written as $F=f_{2}f_{3}f_{1}$, where $f_{j}=\prod_{k=0}^{N/3-1}V_{j+3k,j+3k+1}$ is one layer beginning at site $j$ and $V_{i,j}$ denotes the unitary 2-site gate $V$ acting on qubits $i$ and $j$. Indices are taken modulo $N$, with sites $j=1,\dots,N$.
Figure 1: A chaotic 3-layer BW circuit (inset) and the eigenphases level-spacing distribution $P(s)$. Eigenphases of the propagator $F$ (blue) do not follow the RMT expectation, while after resolving the space-time symmetry the eigenphases of the root $\tilde{F}=(S^{6}F)^{1/3}$ (purple) do agree with the CUE RMT (green curve). Red curve is the theory for a direct sum of 3 CUE matrices (see Eq. (9) and Appendix B). Data is for $N=12$ in the eigenspace with momentum 00, and Haar random gate $V$.
Taking a 2-qubit gate $V$ to be some fixed generic unitary, and therefore having a system that should be quantum chaotic, we can see in Fig. 1 that, after resolving the obvious translational symmetry, the LSD of $F$ is far from the expected RMT result for a circular unitary ensemble (CUE) Haake (2010). If anything, it is closer to a Poisson statistics typical of integrable systems, as if there would be some unresolved symmetry Giraud et al. (2022).
Indeed, each layer $f_{j}$, and thereby also $F$, is invariant under translation by 3 sites, $S^{-3}f_{j}S^{3}=f_{j}$, where $S$ is the translation operator by 1 site to the left,
$V_{i+1,j+1}=S^{-1}V_{i,j}S.$ (1)
We can also easily see (Fig. 1) that translating $F$ by 2 sites is the same as a shift in time by 1 layer. Denoting propagator from time $t_{1}$ to $t_{2}$ by $F(t_{1},t_{2})$, e.g., $F=F(0,1)$, the 3-layer BW circuits has a space-time symmetry $S^{-2}F(0,1)S^{2}=F(1/3,4/3)$ (application of each gate advances time by $1/N$). This symmetry is also reflected in the structure of $F$, which can be written as $F=S^{-4}f_{1}S^{4}S^{-2}f_{1}S^{2}f_{1}=S^{-6}(S^{2}f_{1})^{3}=S^{-6}\tilde{F}^{3}$. We now see where the crux of the problem lies. Since $F$ as well as $f_{1}$ has translational symmetry by 3 sites the momentum $k$ labeling eigenvalues of $S^{3}$ is a good quantum number. In each momentum eigenspace $S^{-6}$ is just an overall phase, and therefore $F$ is, up-to this irrelevant phase, equal to the 3rd power of $\tilde{F}=S^{2}f_{1}$. The quantum chaos conjecture should therefore be applied to $\tilde{F}$ rather than to $F$. Doing that one recovers perfect agreement with the RMT (Fig. 1). It also tells us that, provided $\tilde{F}$ and therefore the circuit is “chaotic”, the LSD of $F$ will be equal to that of the 3rd power of a CUE matrix which is equal to a direct sum of 3 independent CUE matrices, Eq. (9).
The above example is just one possible simple circuit – in our classification it is of type $(q,r)=(3,2)$. We shall classify symmetries of all possible simple quantum circuits, showing that all posses an appropriate space-time symmetry. Space-time symmetries have been discussed before in the solid-state physics context and time-periodic $H(t)$ Xu and Wu (2018); Morimoto et al. (2017); Peng and Refael (2019). An important offshoot will be expressing $F$ essentially as a power of simpler matrix $\tilde{F}$, meaning that in order to probe dynamic (chaotic) properties one needs to study $\tilde{F}$ and not $F$. Expressions of that form have appeared before for special cases of the BW circuit in Ref. Sá et al. (2021) (class $(2,1)$), and for $r=1$ in Ref. Huang et al. (2023), see also Ref. Duh for preliminary results.
Classification of simple circuits.– Having an $N$-site 1D system with periodic boundary conditions there are $N$ nearest-neighbor gates $V_{j,j+1}$ that can be ordered in $N!$ different ways (configurations) to make a one-step propagator $F$  [^2] Single-site Hilbert space can have any dimension (qudits) although we will frequently speak about qubits because all our numerical examples will be of that kind.. However, it is clear that many of those configurations have equal $F$ since 2-site gates acting on non-overlapping nearest neighbor sites commute. Furthermore, a lot of configurations lead to the same spectra, for instance, under cyclic permutations of gates spetra do not change Bensa and Žnidarič (2021). Because we want to study spectra of $F$ we will call two circuits equivalent if they have the same spectrum. It is clear that there are much less than $N!$ non-equivalent simple circuit classes. For open boundary conditions there is in fact just one class Bensa and Žnidarič (2021).
Allowed $(q,r)$
$\mathbf{N}$ generalized S generalized BW
6 (6,1),(6.5) (2,1),(3,1),(3,2)
7 (7,1),(7,2),(7,3),(7,4),(7,5),(7,6)
8 (8.1),(8.3),(8,5),(8,7) (2,1),(4,1),(4.3)
9 (9,1),(9,2),(9,4),(9,5),(9,7),(9,8) (3,1),(3,2)
10 (10,1),(10,3),(10,7),(10,9) (2,1),(5,1),(5,2),(5,3),(5,4)
Table 1: List of $(N-1)$ allowed shift parameters $(q,r)$ for few small $N$, classifying all possible spectrally equivalent $N$ qubit circuits with periodic boundary conditions.
Figure 2: Any same-gate nearest-neighbor circuit is spectrally equivalent to one of the canonical circuits $F_{q,r}$ shown here for $N=6$, see also Table 1.
For periodic boundary conditions this is not the case. One can show (see Appendix, Theorem S1) that there are $(N-1)$ different equivalence classes. A canonical representative circuit of a given equivalence class follows a similar logic as the 3-layer BW example in introduction: a canonical circuit is characterized by two integers $q$ and $r$, where the first layer of gates $f_{1}$ is made by repeatedly translating $V_{1,2}$ by $q$ sites ($q=3$ in the example), whereas $r$ determines the shift of the 2nd layer with respect to the 1st one ($r=2$ in the example). The following theorem embodies the precise statement.
Theorem 1.
Any simple qubit circuit on $N$ sites with periodic boundary conditions is equivalent to exactly one of the $N-1$ canonical simple qubit circuits having Floquet propagator
$F_{q,r}=S^{-qr}\left(S^{r}(S^{q}V_{1,2})^{N/q}\right)^{q},$ (2)
where $q$ is larger than $1$ and divides $N$, and $q$ and $r$ are coprime,
$2\leq q\leq N,$ $\mathrm{gcd}(q,N)=q,$
$1\leq r<q,$ $\mathrm{gcd}(q,r)=1,$ (3)
where $\mathrm{gcd}$ denotes the greatest common divisor.
The proof can be found in Appendix C. While it is not constructive in the sense of providing an explicit procedure of transforming a given $F$ to its canonical form $F_{q,r}$, the transformation is in practice easily achieved by hand for small $N$, or one can in linear time calculate the invariant $p$ introduced in Lemma S1 in Appendix C, thereby obtaining the correct $(q,r)$. The integer invariant $p$ is defined for an alternative simple circuit representative of a given class, and characterizes the circuit as a concatenation of two staircase sections with opposite chirality, the length of the 2nd being $p$.
The canonical form of $F_{q,r}$ in Eq. (2) has a simple geometric interpretation: the term in the inner bracket is a single layer $f_{1}$ that is composed of $N/q$ gates, which is then with appropriate shifts repeated in altogether $q$ layers, explicitly written as
$F_{q,r}=\prod_{j=0}^{q-1}\prod_{i=0}^{N/q-1}V_{1+jr+iq,2+jr+iq}.$ (4)
The $(N-1)$ classes described in Theorem 1 account for all possible circuits of which the standard brick-wall with $(2,1)$, and the staircase Nahum et al. (2018); Lin et al. (2021); Bensa and Žnidarič (2021) (also called convolutional codes Poulin et al. (2009); Anikeeva et al. (2021)) with $(N,1)$ are just two cases. The allowed values of $(q,r)$ for few small $N$ are listed in Table 1, while their pictures are shown for $N=6$ in Fig. 2, and for $N=10$ in Fig. S1 in Appendix. Note that while the allowed set of $(q,r)$ for a given $N$ depends on the factors of $N$, the allowed $N$s for a given $(q,r)$ are simpler: taking any coprime $q>r$ a circuit is possible for all $N$ that are multiples of $q$.
The set of allowed $(q,r)$ naturally splits into two categories. Because $q$ divides $N$, with the maximal value being $N$, one group is composed of the largest possible $q=N$, while the other has smaller $22\leq q\leq N/$ 2. Group (i) are generalized S circuits with $q=N$. Because the translation by $N$ is equivalent to the identity this group could be equivalently described by $(N,r)\equiv(r,0)$, that is, by a single integer $r$ that gives the shift of the next gate. As $r$ is coprime with $N$ all n.n. gates are obtained by a simple (ergodic) translation modulo $N$. Group (ii) can be viewed as generalized BW circuits and needs two integers. Because $q$ divides $N$, translation by $q$ alone is not ergodic and one needs subsequent layers characterized by $r$. Altogether one has a $q$-layer BW circuit, each layer consisting of $N/q$ gates.
Figure 3: Level-spacing distribution $P(s)$ of $F_{q,r}$ (Eq. 2) in blue are not chaotic and are equal to a direct sum of $q$ CUE matrices (red curve, Appendix B), while the LSD of the root $\tilde{F}_{q,r}$ (Eq. 5, violet) agrees with the CUE RMT prediction (green curve). For $q\neq N$ we show data from the momentum eigenspace of $S^{q}$ with $k=0$, and averaging is done over $10-50$ circuits.
For a generic gate $V$ the spectra of all $(N-1)$ propagators $F_{q,r}$ are different. If $V$ would be symmetric with respect to the exchange of the two qubits the circuits with $(q,r)$ and $(q,q-r)$ would have the same spectra (spatial reflection symmetry), and therefore one would have only $\lfloor\frac{N}{2}\rfloor$ different spectral classes Bensa and Žnidarič (2021). Theorem 1 also shows that for prime $N$ only generalized S circuits exist. For odd $N$ there are no standard BW circuits having $(2,1)$, but there are generalized BW circuits with $q>2$ (see Table 1). Note also that while for smaller $N<10$ possible circuits are straightforward generalizations of the S or BW configurations with left or right chirality, for larger $N$ less intuitive circuits are also possible. For instance, for $N=10$ one can have $(5,2)$ (see Fig. S1 in Appendix) that can be further compressed in time direction (e.g., all gates in the first two layers $f_{1}$ and $f_{3}$ commute), reducing the number of non-commuting layers from $q=5$ to just $3$. Each of those compressed layers has $2$ idle qubits (that are not acted upon), such that the compressed circuit $F_{5,2}^{t}$ has two separate lines of idle qubits, each of width $1$, slanted at $-45^{\circ}$. Integers $(q,r)$ therefore also determine the filling fraction, i.e., the number and the pattern of idle qubits in maximally compressed $F_{q,r}^{t}$ (see Appendix A). The only circuit with no idle qubits is the standard BW with $(2,1)$.
Space-time symmetries.– In order to understand space-time symmetries of any simple circuit it suffices to study the canonical equivalence class representatives $F_{q,r}$. Denoting the inner term in Eq. (2) by $\tilde{F}_{q,r}$, calling it a root of $F_{q,r}$,
$\tilde{F}_{q,r}$ $=S^{r}(S^{q}V_{1,2})^{N/q}=S^{r}f_{1},$ (5)
where $f_{1}=\prod_{k=0}^{N/q-1}S^{-kq}V_{1,2}S^{kq}$, we can write Eq. (2) as
$F_{q,r}=S^{-qr}(\tilde{F}_{q,r})^{q}.$ (6)
Eq. (6) appeared in Ref. Sá et al. (2021) in the open-systems context for the special case of a BW circuit with $(2,1)$. The root connection is especially simple for the generalized S case: for $q=N$ the translation $S^{q}$ is identity, resulting in $\tilde{F}_{q,r}=S^{r}V_{1,2}$ and $F_{q,r}=(S^{r}V_{1,2})^{N}$.
The root has a translational symmetry by $q$ sites
$S^{-q}$ $\tilde{F}_{q,r}S^{q}=\tilde{F}_{q,r}.$ (7)
This is trivially true if $q=N$, otherwise the translated root is equal to $S^{r}\prod_{k=0}^{N/q-1}S^{-(k+1)q}V_{1,2}S^{(k+1)q}$, where, since all the gates in the product commute, we can relabel the index $k+1=k^{\prime}$, yielding $S^{r}\prod_{k^{\prime}=0}^{N/q-1}S^{-k^{\prime}q}V_{1,2}S^{k^{\prime}q}=\tilde{F}_{q,r}$. Because $F_{q,r}$ is a power of $\tilde{F}_{q,r}$ multiplied by some power of $S^{q}$, $F_{q,r}$ also has translational symmetry by $q$ sites.
Furthermore, $F_{q,r}$ also has a space-time symmetry when translating by $r$ sites
$S^{-r}F_{q,r}(0,1)S^{r}=F_{q,r}(1/q,1+1/q).$ (8)
This can be easily seen by rewriting the LHS of Eq. (8) as $S^{-qr}\left(S^{r}S^{-r}(S^{q}V_{1,2})^{N/q}S^{r}\right)^{q}$, which then equals to $S^{-qr}\left(S^{r}(S^{q}V_{1+r,2+r})^{N/q}\right)^{q}$. The final expression can be understood as a circuit beginning with the second layer, thus justifying the equality to the RHS of Eq. (8). Eq. (8) appeared in Ref. Huang et al. (2023) for the special case of $r=1$.
Level spacing statistics.– An immediate application of the above results is in quantum chaos for the statistics of spacings of closest eigenphases of $F$. Looking at the root connection in Eq. (6) and the fact that $S^{q}$ commutes with all terms, one can focus on a given common momentum eigenspace of $S^{q}$ with eigenvalues $e^{2\pi{\rm i}kq/N}$, $k\in\{0,1,\dots,N/q-1\}$. There $S^{-qr}$ is just an overall phase factor $e^{-2\pi{\rm i}qkr/N}$. Therefore $F_{q,r}$ is up-to this phase equal to an appropriate power of $\tilde{F}_{q,r}$. Eigenphases of $F_{q,r}$ are simple multiples of eigenphases of $\tilde{F}_{q,r}$ modulo $2\pi$, and therefore when trying to test the validity of the quantum chaos conjecture one must consider the level-spacing distribution of the root $\tilde{F}_{q,r}$ instead of the whole Floquet propagator $F_{q,r}$.
If the circuit is quantum chaotic one would therefore expect that the spectral statistics of the root $\tilde{F}_{q,r}$ (5) will adhere to the RMT theory. In particular, if the 2-site gate $V$ does not have any anti-unitary (time-reversal) symmetry, which is the case for our numerics where $V$ is randomly picked according to the Haar measure, the appropriate ensemble is the CUE. We can see in Fig. 3 that the LSD of $\tilde{F}_{q,r}$ indeed agrees with CUE Wigner surmise $P(s)=\frac{32}{\pi^{2}}s^{2}e^{-s^{2}4/\pi}$ for all canonical classes. Furthermore, the theorem by Rains Rains (2003) tells us what is the distribution of a power of a matrix from the CUE. For $M\in\mathrm{CUE}(\mathcal{N})$, where $\mathrm{CUE}(\mathcal{N})$ denotes a distribution of $\mathcal{N}\times\mathcal{N}$ CUE matrices, one has
$M^{q}\in\bigoplus_{j=1}^{q}\mathrm{CUE}\left(\left\lceil\frac{\mathcal{N}-j}{q}\right\rceil\right),$ (9)
[…]
Discussion.– We have classified all different quantum circuits in one dimension with periodic boundary conditions and translationally invariant nearest-neighbor 2-site gates. There are $N-1$ different spectral classes, being generalizations of the familiar brick-wall and the staircase configurations. Each class can be characterized by two integers $(q,r)$, such that the Floquet propagator is essentially a $q$-th power of $\tilde{F}=S^{r}f_{1}$, where for generalized S circuits one has $f_{1}=V_{1,2}$, while for generalized BW circuits $f_{1}$ is one layer of gates. We have therefore come full circle: similarly as in classical single-particle kicked models where one interchangeably applies a simple map in real space (e.g., potential $V(q)$) and a simple map in momentum space (e.g., free evolution), any quantum many-body translationally invariant Floquet system has the same basic structure. The elementary building block $\tilde{F}$ is a product of simple local transformation, like $V_{1,2}$, and of “free evolution” described by the translation operator $S$ (that is diagonal in the Fourier basis).
We have explicitly shown how that affects the level spacing statistics of simple circuit Floquet systems – to detect quantum chaos one must look at the spectrum of the $q$-th root $\tilde{F}$ of the propagator. Effects of the underlying space-time symmetry on other quantifiers of quantum chaos remain to be explored. Reducing all circuits to just few canonical classes makes it possible to study chaoticity for different $(q,r)$; are some configurations more chaotic than others, does that depend on the filling fraction (see Appendix A)? While we focused on systems without any symmetry, i.e., the unitary case, orthogonal and symplectic cases can be treated along the same lines. Generalizing classification to more than one dimension is also an open problem.
Acknowledgements.– We acknowledge support by Grants No. J1-4385 and No. P1-0402 from the Slovenian Research Agency.
Appendix A Canonical circuits
In the main text in Fig. 2 we have shown canonical circuits $F_{q,r}$ for $N=6$. Here we show in Fig. S1 all $9$ canonical $F_{q,r}$ for $N=10$, where a variety of configurations is richer. We can for instance notice that even though circuits are constructed as a $q$-layered circuit in some cases consecutive layers commute and can therefore be compressed, thus reducing the number of layers. For $N=10$ this is the case for $(q,r)=(5,2)$ and its chiral pair $(5,3)$. For $(5,2)$ the first two layers $f_{1}$ and $f_{3}$ commute and can be compressed to a single layer; likewise for the next two layers. Therefore in the compressed form $(F_{5,2})^{2}$ consists of only 5 layers instead of $10$. In this compressed form there are still $2$ idle qubits in each layer on which no gate acts. One can say that the filling fraction of gates for the circuit $(5,2)$ is $8/10$, i.e., $20\%$ of qubits is idle. The class $(5,1)$ on the other hand can not be compressed any further and has the filling fraction of only $4/10$. The only circuit with filling fraction $1$ is the standard BW with $(2,1)$.
Appendix B Level-spacing distribution of a direct sum of independent RMT matrices
The general formula for the LSD of a direct sum of independent RMT matrices of arbitrary dimensions was derived in Ref. Rosenzweig and Porter (1960). The matrices in the direct sum in the theorem by Rains (Eq. (9), Ref. Rains (2003)) are approximately of equal dimensions in the large $N$ limit, which is why we use the reduced formula for the direct sum of equally dimensional matrices throughout this paper.
For the LSD $P(s)$ of an RMT ensemble we define
$R(y)$ $=\int_{0}^{\infty}P(x+y)\,\mathrm{d}x,$ (10)
$D(y)$ $=\int_{0}^{\infty}xP(x+y)\,\mathrm{d}x,$ (11)
where $R(y)$ is nothing but $1$ minus the cumulative distribution of $P$. The LSD for a direct sum of $m$ independent equally dimensional matrices from the same RMT ensemble is then
$P_{m}(s)=D^{m}(s/m)\left[\frac{1}{m}\frac{P(s/m)}{D(s/m)}+\left(1-\frac{1}{m}\right)\frac{R^{2}(s/m)}{D^{2}(s/m)}\right].$ (12)
In Figures in the paper, we plot $P_{m}(s)$ obtained by using the Wigner surmise for $P(s)$, which is of acceptable accuracy for our application. For better approximations of $P(s)$ see Ref. Dietz and Haake (1990).
Appendix C Proof of Theorem 1 from the main paper
In this section, we set to prove Theorem 1, the main result of this paper. The proof is divided into subsections containing more theorems and lemmas. First, we introduce another canonical configuration in Section C.1, which can be geometrically interpreted as a concatenation of two staircase circuits with opposite chiralities. It is useful for determining that the number of non-equivalent circuits is $(N-1)$ and later used in the proof of Theorem 1. After that in Section C.2, we introduce the quantity $C(F)$ invariant for equivalent circuits and state its important properties. This invariant is crucial in later proofs but is also useful on its own to efficiently determine the canonical form of a given circuit. Finally, we focus on the generalized S/BW circuits $F_{q,r}$ in Section C.3, where we combine the results from previous sections to prove Theorem 1.
To keep our notation clearer, we identify a product of $N$ 2-site gates acting on nearest neighbor sites with a sequence of $N$ numbers in the following way
$V_{i_{N},i_{N}+1}\cdots V_{i_{2},i_{2}+1}V_{i_{1},i_{1}+1}\equiv(i_{1},i_{2},\dots,i_{N}).$ (13)
We will refer to $i_{l}$ as gate numbers and $l$ as time indices. When talking about the gates appearing before/after a gate $i_{l}$ in a given $F$, we will call $i_{l+1}$ the time successor and $i_{l-1}$ the time predecessor, whereas when refering to gates with neighboring numbers, we will call $i_{l}-1$ the left neighbor and $i_{l}+1$ the right neighbor of $i_{l}$. By definition, the sequences corresponding to simple circuits are permutations of the first $N$ natural numbers. For the canonical simple circuits defined in the paper in Eq. 2
$F_{q,r}\equiv($ $1,1+q,\dots,1+(N/q-1)q,$
$1+r,1+r+q,\dots),$ (14)
where all gate numbers are taken modulo $N$ (from $1$ to $N$).
We are interested in (spectrally) equivalent circuits, as defined in the paper. The equivalence will be denoted with $\cong$. The two important equivalence operations are time predecessor/successor commutation of non-neighboring gates (here the Floquet operators are actually equal, not only equivalent)
$|i_{k}-i_{k+1}|$ $>1\pmod{N}\quad\implies$
$V_{i_{N},i_{N}+1}$ $\cdots V_{i_{k+1},i_{k+1}+1}V_{i_{k},i_{k}+1}\cdots V_{i_{1},i_{1}+1}=$ (15)
$=V_{i_{N},i_{N}+1}\cdots V_{i_{k},i_{k}+1}V_{i_{k+1},i_{k+1}+1}\cdots V_{i_{1},i_{1}+1},$
in the new notation written as
$|i_{k}-i_{k+1}|>1\pmod{N}\quad\implies$ (16)
$\ (i_{1},\dots,i_{k},i_{k+1},\dots,i_{N})\cong(i_{1},\dots,i_{k+1},i_{k},\dots,i_{N}),$
and cyclic permutation
$V_{i_{N},i_{N}+1}\cdots$ $V_{i_{2},i_{2}+1}V_{i_{1},i_{1}+1}\cong$
$\cong V_{i_{1},i_{1}+1}V_{i_{N},i_{N}+1}\cdots V_{i_{2},i_{2}+1},$ (17)
or in the new notation
$(i_{1},i_{2},\dots i_{N})\cong(i_{2},\dots,i_{N},i_{1}).$ (18)
While the equivalence of cyclically permuted circuits was motivated by spectral equivalence, a different, perhaps more general, motivation is also possible. If we have dynamics in mind, i.e., $F^{t}$ with possibly large $t$, the definition of a starting time of our period is arbitrary, e.g., the first operator in a period could just as well have been defined as the last operator in the previous period. Therefore, it would also makes sense to define cyclically permuted circuits to be equivalent in this case.
C.1 Double staircase canonical circuits $F_{p}$
We now introduce a new canonical form, different from the one in the main paper, its diagram is shown in Fig. S2. It consists of two staircase sections with opposite chirality, the second having length $p$. To show that it is indeed a canonical form, i. e. that every simple circuits is equivalent to some circuit in this form, we can prove the following Theorem.
Figure S2: Diagram of $F_{p}$ defined in Eq. (19).
Theorem S1.
Any simple circuit is equivalent to a simple circuit with the Floquet operator given by
$F_{p}=$ $V_{N-p+1,N-p+2}\cdots V_{N-1,N}V_{N,1}$
$V_{N-p,N-p+1}\cdots V_{2,3}V_{1,2}$ (19)
[…]
for some $p\in\{1,2,\dots,N-1\}$.
Proof.
Let $F=(i_{1}^{(0)},i_{2}^{(0)},\dots,i_{N}^{(0)})$ be any simple circuit. This means that $i_{k}^{(0)}\neq i_{l}^{(0)}$ for $k\neq l$. First, we will bring gate $1$ to index $1$ (step 1). After that, we will try to bring gate $2$ to index $2$ (step 2). This will either be possible, in which case we will continue to try bring gate $k$ to index $k$ in the general step, or it won’t be possible, in which case our gate will be equivalent to $F_{N-1}$.
Step 1: By means of cyclic permutations (18) we can always transform $F$ into an equivalent simple circuit beginning with gate $1$ ($V_{1,2}$ in the standard notation)
$F\cong(1,i^{(1)}_{2},\dots,i^{(1)}_{N}),$ (20)
where we have appropriately relabeled the indices.
Step 2: We now try to bring gate $2=i^{(1)}_{K},K\in\{2,\dots N\}$ to the second position by doing time predecessor/successor commutations (16). Gate $2$ does not commute only with gates $1$ and $3$, which means that we must consider two possibilities:
  1. (i)
    Gate $3$ appears after gate $2$, $3=i^{(1)}_{L}$ for $L>K$:
    By applying (16), we can bring gate $2$ to index $2$
    $F\cong(1,2,i^{(2)}_{3},\dots,i^{(2)}_{N}),$ (21)
    where we have again relabelled the indices. We can continue with the general step.
  2. (ii)
    Gate $3$ appears before gate $2$, $3=i^{(1)}_{L}$ for $L<K$:
    By applying (16) (and relabelling the indices), we can bring gate $2$ to be the time successor of gate $3$
    $F\cong(1,i^{\prime(1)}_{2},\dots,3,2,\dots,i^{\prime(1)}_{N}).$ (22)
    In the context of equivalence transformations, we can now think of the sequence of gates $(3,2)=(i^{\prime(1)}_{K^{\prime}},i^{\prime(1)}_{K^{\prime}+1})$ as a gate that does not commute only with gates $1$ and $4$. Again, we have two similar cases:
    1. (a)
      Gate $4$ appears after the sequence of gates $(3,2)$, $4=i^{\prime(1)}_{L^{\prime}}$ for $L^{\prime}>K^{\prime}+1$:
      By applying (16), we can bring gate $3$ to index $1$ and gate $2$ to index $3$
      $F\cong(3,1,2,\dots)$ (23)
      and by applying (18), we can cyclically permute gate $3$ to index $N$
      $F\cong(1,2,i^{(2)}_{3},\dots,i^{(2)}_{N-1},3).$ (24)
      We can now continue with the general step.
    2. (b)
      Gate $4$ appears before $(3,2)$, $4=i^{\prime(1)}_{L^{\prime}}$ for $L^{\prime}<K^{\prime}$:
      By applying (16)
      $F\cong(1,\dots,4,3,2,\dots).$ (25)
      We can again think of $(4,3,2)$ as a gate that does not commute only with gates $5$ and $1$ and again consider two cases similar to (a) and (b). By repeating this process, we either get to the point where case (a) arises, and we can continue with the general step, or our circuit is equivalent to
      $F\cong(1,N,N-1,\dots,2).$ (26)
      Here, the obtained equivalent circuit is thus precisely $F_{N-1}$.
General step: Let’s suppose we have already shown that $F$ is equivalent to
$F\cong(1,2,\dots,k,i^{(k)}_{k+1},\dots,i^{(k)}_{N}),$ (27)
where $i^{(k)}_{l}$ are arbitrary indices relabelled in a convenient way. We can treat the sequence of gates $(1,2,\dots,k)$ as a gate, that does not commute only with $1$ and $k+1$, which means that we can repeat step 2 by trying to bring gate $k+1$ to the right of $k$ (index $k+1$). Thus, either
  1. (i)
    $F\cong(1,\dots,k,k+1,\dots)$ or
  2. (ii)
    $F\cong(1,\dots,k,N,N-1,\dots,k+1)=F_{N-k}$.
In (i) we can repeat the general step until case (ii) arises, or we end up with $F\cong(1,2,\dots,N)=F_{1}$. ∎
Since there are $(N-1)$ allowed $p$, it is clear that there are (at most) $(N-1)$ non-equivalent simple circuits. The proof given is constructive, which means that we can use it as an algorithm to convert a given $F$ to some $F_{p}$.
C.2 Circuit invariant
An alternative way to determine the equivalent $F_{p}$ for a given $F$ is to calculate some quantity, which is invariant in equivalence operations and is different for all $F_{p}$. A convenient choice is the length of the second staircase $p$, which is clearly equal to the number of gates, for which their right neighbor (modulo $N$) appears in the Floquet operator before them. Let’s thus define $C(F)$ to be exactly that
[…]
$i_{k}-i_{l}\equiv 1\pmod{N}\}|,$ (28)
where ||$\bullet$|| ∙ | denotes the cardinality of a set. As stated, $C(F_{p})$ is clearly equal to the length of the second staircase
$C(F_{p})=|\{N,N-1,\dots,N-p+1\}|=p.$ (29)
Let’s now show that $C(F)$ is indeed invariant under circuit equivalence operations.
Lemma S1.
For simple circuits $C(F)$ is invariant under circuit equivalence operations (16) and (18).
Proof.
Let
[…]
$i_{k}-i_{l}\equiv 1\pmod{N}\},$ (30)
be the set of gates the invariant is counting, which means $C(i_{1},\dots,i_{N})=|\tilde{C}(i_{1},\dots i_{N})|$.
As stated previously, for simple circuits $(i_{1},\dots,i_{N})$ is just a permutation of $(1,\dots,N)$, which means
$i_{1}$ $\notin\tilde{C}(i_{1},\dots,i_{N}),\text{because nothing is before}\ i_{1}$ (31)
$i_{1}-1$ $\in\tilde{C}(i_{1},\dots,i_{N}),\text{because}\ i_{1}-1\in\{i_{j}\}_{j=2}^{N},$ (32)
where indices are taken modulo $N$. But in the equivalent cyclically permuted circuit (18)
$i_{1}$ $\in\tilde{C}(i_{2},\dots,i_{N},i_{1}),\text{because}\ i_{1}+1\in\{i_{j}\}_{j=2}^{N}$ (33)
$i_{1}-1$ $\notin\tilde{C}(i_{2},\dots,i_{N},i_{1}).$ (34)
The membership of all other gates in $\tilde{C}$ stays the same in both cases, since the position of $i_{1}$ can change only the membership of $i_{1}-1$ and $i_{1}$ in $\tilde{C}$ and all other relative positions stay the same. Thus
$C(i_{1},\dots,i_{N})=C(i_{2},\dots,i_{N},i_{1}),$ (35)
which means that $C$ is conserved under cyclical permutations (18).
A transposition of time successive gates can only change $C$ if their gate number difference is ±$\pm 1$, which does not happen if the commute (Eq. (16)). Therefore, $C$ is conserved there
$|i_{k}-i_{k+1}|>1\pmod{N}\quad\implies$
$\ C(i_{1},\dots,i_{k},i_{k+1},\dots,i_{N})=C(i_{1},\dots,i_{k+1},i_{k},\dots,i_{N}).$ (36)
An obvious consequence of Theorem S1 is the following lemma.
Lemma S2.
For simple circuits $C(F)\in\{1,\dots,N-1\}$.
Proof.
According to Theorem S1, any simple circuit $F$ is equivalent to some $F_{p}$ for $p\in\{1,\dots,N-1\}$. Since $C(F)$ is conserved under equivalence transformations (Lemma S1) and $C(F_{p})=p$, $C(F)\in\{1,\dots,N-1\}$. ∎
C.3 Generalized S/BW canonical circuits $F_{q,r}$
We now turn to the canonical form $F_{q,r}$. For a given $N\in\mathbb{N}$, define the set of allowed $(q,r)$ pairs (see Theorem 1)
$Q_{N}=\{(q,r);$ $2\leq q\leq N,1\leq r<q,$
$\mathrm{gcd}(q,N)=q,\mathrm{gcd}(q,r)=1\}.$ (37)
The allowed $(q,r)$ generate either a generalized S circuit
$Q_{N}^{(\mathrm{S})}=\{(N,r);1\leq r<N,\mathrm{gcd}(N,r)=1\}$ (38)
or a generalized BW circuit
$Q_{N}^{(\mathrm{BW})}=\{(q,r);$ $2\leq q<N,1\leq r<q,$
$\mathrm{gcd}(q,N)=q,\mathrm{gcd}(q,r)=1\}.$ (39)
Which means that $Q_{N}=Q_{N}^{(\mathrm{S})}\cup Q_{N}^{(\mathrm{BW})}$, where $Q_{N}^{(\mathrm{S})}\cap Q_{N}^{(\mathrm{BW})}=\emptyset$.
The first thing to check is that $F_{q,r}$ is well defined, i. e. if Floquet operators $F_{q,r}$ actually correspond to simple circuits.
Lemma S3.
$F_{q,r}$ for $N\in\mathbb{N}$ and $(q,r)\in Q_{N}$ are simple circuits, i. e. contain every gate on neighboring sites exactly once.
Proof.
To show that $F_{q,r}$ are simple circuits, we thus have to check that the sequence (14) contains every positive integer less than or equal to $N$ exactly once.
Let’s first consider the generalized S case, where $q=N$ and $\mathrm{gcd}(N,r)=1$. Because of that, it is intuitively clear, that translations by $r$ must be “ergodic” and thus generate all gates. More formally, this means that the least common multiple $\mathrm{lcm}(N,r)=rN/\mathrm{gcd}(N,r)=rN$. Since $F_{N,r}$ is generated by translating gate $1$ by $r$ sites, the lowest number of translations after which gate $1$ (modulo $N$) is generated again must be $N$ (since the smallest solution to $kr=0\pmod{N}$ is exactly $k=\mathrm{lcm}(N,r)/r$). Thus, before looping back to gate $1$ we generate $N$ gates. By an analogous argument, any gate is generated again only after $N$ translations, which means that all the generated gates are different.
Let’s now consider the generalized BW case where $q\neq N$. In this case $\mathrm{lcm}(q,N)=N$, which means that with translations by $q$ we generate $N/q$ different gates. These are exactly gates $1$ plus all the multiples of $q$ less than or equal to $N$, so gates $1+qk$, $k=0,\dots,N/q-1$. By translating them by any $r$, such that $1\leq r<q$, we generate all gates $1+r+qk$, which since $r<q$ cannot be equal to any gates generated with $r=0$. More formally, this is true because $1+r+qk\not\equiv 1+qk\pmod{q}$ necessarily implies $1+r+qk\neq 1+qk$. In order to generate all gates, $1+kr$ for $k\in\{0,1,2,\dots,q-1\}$ must now take all the possible values modulo $q$. Analogous to the generalized S case, this indeed does happen when $\mathrm{gcd}(q,r)=1$. ∎
Since Theorem S1 implies that there are $(N-1)$ non-equivalent simple circuits, a required condition for $F_{q,r}$ to be a canonical form is that there are $(N-1)$ allowed $(q,r)$.
Lemma S4.
$|Q_{N}|=N-1$.
Proof.
Let $\varphi(N)=|\{k;1\leq k<N,\mathrm{gcd}(N,k)=1\}|$ denote the number of natural numbers less than $N$ and coprime with $N$, also called the Euler’s totient or Euler’s phi function.
We can now express the number of generalized staircase circuit with Euler’s phi function
$\left|Q_{N}^{(\mathrm{S})}\right|=|\{(N,r);1\leq r<N,\mathrm{gcd}(N,r)=1\}|=\varphi(N)$ (40)
and also the number of generalized BW circuits
[…]
$\quad\mathrm{gcd}(q,N)=q,\mathrm{gcd}(q,r)=1\}|$
$=$ $\ \sum_{q=2,q|N}^{N-1}|\{(q,r);1\leq r<q,\mathrm{gcd}(q,r)=1\}|$
$=$ $\ \sum_{q=2,q|N}^{N-1}\varphi(q)$ (41)
where $q|N$ denotes that $q$ divides $N$ (equivalently $\mathrm{gcd}(q,N)=q$).
Thus, since $Q_{N}^{(\mathrm{S})}\cap Q_{N}^{(\mathrm{BW})}=\emptyset$
$|Q_{N}|$ $=\left|Q_{N}^{(\mathrm{S})}\right|+\left|Q_{N}^{(\mathrm{BW})}\right|=\sum_{q=2,q|N}^{N}\varphi(q)$
$=\sum_{q|N}\varphi(q)-\varphi(1)=N-1,$ (42)
where we used the well-known theorem that $\sum_{q|N}\varphi(q)=N$ Rosen (2011) and $\varphi(1)=1$. ∎
We now wish to show that the invariant $C(F_{q,r})$ is different for all allowed $(q,r)$, which is the last important statement required for the proof of Theorem 1. We do this in two steps, first for the generalized S in Lemma S5 and then for generalized BW in Lemma S6, finally combining both in Lemma S7.
Lemma S5.
In the generalized S case, $(q,r)\in Q_{N}^{(\mathrm{S})}$,
$C(F_{q,r})$ $=1+|\{\text{gates appearing after gate $N$}\}|$ (43)
$\in\{p;p<N,\mathrm{gcd}(p,N)=1\},$ (44)
and is different for different $r$. Here all the values in the set (44) are taken for some allowed $r$.
Proof.
Since we are considering the generalized S case: $q=N$, $r<N$ and $\mathrm{gcd}(N,r)=1$.
We first want to consider if we know that a certain gate is a member of $\tilde{C}(F_{N,r})$, what can we say about the membership of its time successor and predecessor. Let $F_{N,r}=(i_{1},\dots,i_{N})$. Let $i_{l}\in\tilde{C}(F_{N,r})$, which means that its right neighbor appears in $F_{N,r}$ before it, $\exists i_{k}=i_{l}+1,k<l$ (gate numbers are taken modulo $N$, but indices are not). Here $l\neq 1$, since in that case, such $k<l=1$ never exists. Let’s now consider the membership of $i_{l}$’s successor and predecessor:
  1. (a)
    If $i_{l+1}=i_{l}+r$ is a valid gate (i. e. its index is valid, which means $l<N$), then $i_{k+1}=i_{k}+r=i_{l}+1+r=i_{l+1}+1$, is the right neighbor of $i_{l+1}$ and $k+1<l+1$, which implies $i_{l+1}\in\tilde{C}(F_{N,r})$. In other words: If $i_{l}$’s time successor exists (i. e. $i_{l}$ is not the last gate in $F$), it is also a member of $\tilde{C}(F_{N,r})$.
  2. (b)
    $i_{l-1}=i_{l}-r$, then if $i_{k-1}=i_{k}-r=i_{l}+1-r=i_{l-1}+1$ is a valid gate (i. e. $k>1$), it is the right neighbor of $i_{l-1}$ and $k-1<l-1$, so clearly $i_{l-1}\in\tilde{C}(F_{N,r})$. In other words: If $i_{l}$’s right neighbor is not $i_{1}$, then it’s time predecessor is also a member of $\tilde{C}(F_{N,r})$.
In case (b) $i_{k-1}$ is not valid only if $k=1$. In the case of $F_{q,r}$, $i_{1}=1$ and thus $i_{l}=N$. Since always $N\in\tilde{C}(F_{N,r})$ (its right neighbor is $1=i_{1}$), by iterating (a), all gates appearing after gate $N$ are also members of $\tilde{C}(F_{N,r})$. If any gate appearing before gate $N$ would be a member of $\tilde{C}(F_{N,r})$, iterating (b) would eventually lead to $i_{1}=1\in\tilde{C}(i_{1},\dots,i_{N})$, which cannot happen, thus leading to a contradiction. We have shown:
$C(F_{N,r})$ $=1+|\{\text{gates appearing after gate $N$}\}|=$
$=N-k_{N}+1,$ (45)
where $k_{N}$ is the index of gate $N$, $i_{k_{N}}=N$. We have thus shown the first part of the lemma.
We now want to determine what are the possible values of $C(F_{N,r})$. In order to do that, we must only find the possible values of $k_{N}$. By definition, we get gate $N$ after $k_{N}-1$ translations of gate $1$ by $r$
$1+(k_{N}-1)r$ $\equiv 0\pmod{N},$ (46)
$\implies-(k_{N}-1)r$ $\equiv 1\pmod{N}.$ (47)
Thus, $-r$ is a modular multiplicative inverse of $k_{N}-1$. According to a well known theorem Rosen (2011), a requirement for modular multiplicative inverses to exists is $\mathrm{gcd}(k_{N}-1,N)=1$. Therefore
$\mathrm{gcd}(C(F_{N,r}),N)$ $=\mathrm{gcd}(N-k_{N}+1,N)=$
$=\mathrm{gcd}(k_{N}-1,N)=1.$ (48)
In other words, $C(F_{N,r})$ is coprime with $N$. According to Lemma S2, $C(F)\leq N-1$, which means that $C(F_{N,r})$ can only be a number less than $N$ and coprime with $N$
$C(F_{N,r})\in\{p;p<N,\mathrm{gcd}(p,N)=1\}.$ (49)
Let’s now show that $k_{N}$ is different for different $F_{N,r}$ and $F_{N,\tilde{r}}$ by contradiction. Let’s suppose otherwise, this means that $k_{N}-1$ translations by $r$ and by $\tilde{r}$ must generate the same gate number modulo $N$.
$(k_{N}-1)r\equiv(k_{N}-1)\tilde{r}\pmod{N}.$ (50)
Since $\mathrm{gcd}(k_{N}-1,N)=1$, we can divide the equation by $k_{N}-1$ Rosen (2011)
$r\equiv\tilde{r}\pmod{N}.$ (51)
By definition of allowed $F_{N,r}$ circuits, $r,\tilde{r}<N$, which means that $r=\tilde{r}$, leading to a contradiction.
We have thus shown that for different allowed $r$, $C(F_{N,r})$ are different. Moreover, in Eq. (49) we have shown that the value of $C(F_{N,r})$ is a member of a set with exactly the same number of elements as allowed $r$ (see $\left|Q_{N}^{(\mathrm{S})}\right|$ in the proof of Lemma S4). This means that $C(F_{N,r})$ takes all the values from the set in Eq. (49) if we let $r$ be all the allowed $r$ values. ∎
Lemma S6.
In the generalized BW case, $(q,r)\in Q_{N}^{(\mathrm{BW})}$, for fixed $q$
$C(F_{q,r})$$\in\left\{\frac{N}{q}p;p<q,\mathrm{gcd}(p,q)=1\right\}.$ (52)
The values are different for different $r$ and are all taken for some allowed $r$.
Proof.
Since we are considering the generalized BW case: $q|N,r<q,\mathrm{gcd}(q,r)=1$.
In this case, we can divide the circuit in $q$ blocks (layers) of $N/q$ time consecutive gates. The first block is generated by translating gate $1$ by $q$ (modulo $N$) until we reach gate $1$ again. The second block consists of all the gates from the first block translated by $r$. This means that the first block can be mapped to the subgroup of the cyclic group $\mathbb{Z}_{N}=\left\{0,\dots,N-1\right\}$ of order $N/q$, where group addition is taken modulo $N$. The other blocks are cosets of this subgroup (since according to Lemma S3, they must be disjoint), implying that the first $q$ gate numbers are contained in different blocks. We can therefore denote the blocks with the minimal gate number contained in it (and a tilde for clarity) $\tilde{1},\tilde{2},\dots,\tilde{q}$.
We now wish to consider a similar thing as in the proof of Lemma S5, but in the context of a block, i. e. given that a gate is a member of $\tilde{C}$, what can we say about the membership of its time successor and predecessor. We will see that the main difference is that this time, the notion time successor/predecessor can be taken inside a given block, so with indices modulo $N/q$.
Let’s suppose that a gate with index $l$ (this index is now taken in the context of a block, meaning modulo $N/q$) in block $\tilde{l}$ (modulo $q$) is a member of $\tilde{C}(F_{q,r})$, $i_{l}^{\left(\tilde{l}\right)}\in\tilde{C}(F_{q,r})$. This means that its right neighbor must appear before it, $\exists i_{k}^{\left(\tilde{k}\right)}=i_{l}^{\left(\tilde{l}\right)}+1,\tilde{k}<\tilde{l}$ (if such $k$ would exists in block $\tilde{k}=\tilde{l}$, the block would contain all $N$ gate numbers, which is not possible in the generalized BW case). Analogously to the generalized S case, $i_{l+1}^{\left(\tilde{l}\right)}=i_{l}^{\left(\tilde{l}\right)}+q$ and $i_{k+1}^{\left(\tilde{k}\right)}=i_{k}^{\left(\tilde{k}\right)}+q=i_{l}^{\left(\tilde{l}\right)}+q+1=i_{l+1}^{\left(\tilde{l}\right)}+1$ and thus $i_{l+1}^{\left(\tilde{l}\right)}\in\tilde{C}(F_{q,r})$. Crucially, indices $k$ and $l$ here are taken modulo $N/q$ because translating the last gate of the block by $q$ (modulo $N$) yields the first element of the block. Iterating this, we can conclude that if a gate from a block is contained in $\tilde{C}(F_{q,r})$, then all the gates from this block are contained in $\tilde{C}(F_{q,r})$.
We have therefore shown that by relabelling the considered circuit with $\tilde{1},\tilde{2},\dots,\tilde{q}$, the considered circuit behaves exactly the same as is in the generalized S case. Using Lemma S5, for a fixed $q$ and $N$
$C(F_{q,r})\in\left\{\frac{N}{q}p;p<q,\mathrm{gcd}(p,q)=1\right\},$ (53)
where we took into the account that the number of gates in a block is $N/q$. Also according to Lemma S5, all the values from the set in Eq. (53) are taken for some $r$ and are different for different $r$. ∎
Lemma S7.
$C(F_{q,r})$ is different for all $(q,r)\in Q_{N}$ for a given $N$.
Proof.
We have calculated all the allowed values of $C(F_{q,r})$ and shown that they are different for different $r$ and a fixed $q$ in Lemma S5 and Lemma S6. The only thing left to do is to show that for different $q$, the sets (44) and (52) are disjoint.
From Eq. (52), we can see that in the generalized BW case $\mathrm{gcd}(C(F_{q,r}),N)\geq N/q$, meaning that all the values of $C(F_{q,r})$ must be different than in the generalized S case ($q=N$) where $\mathrm{gcd}(C(F_{N,r}),N)=1$ (Eq. (44)).
The only thing left is to check that the sets in the BW case (Eq. (52)) are disjoint for two different $q_{1},q_{2}|N$. We can do this by finding a contradiction in the converse case. Let’s suppose that we would find the same value of $C$ for different $q_{1},q_{2}$, this means that for some $p_{1},p_{2},\mathrm{gcd}(p_{1},q_{1})=\mathrm{gcd}(p_{2},q_{2})=1$ we must have
$\frac{N}{q_{1}}p_{1}$ $=\frac{N}{q_{2}}p_{2},$ (54)
$\implies q_{2}p_{1}$ $=q_{1}p_{2}.$ (55)
This implies that $p_{1}|q_{1}p_{2}$, but since $\mathrm{gcd}(p_{1},q_{1})=1$, we see that $p_{1}|p_{2}$. Symmetrically, we can show $p_{2}|p_{1}$, which then implies $p_{1}=p_{2}$. From Eq. (55) we finally get $q_{1}=q_{2}$, a contradiction. Therefore, $C(F_{q,r})$ must also be different for different $q$ in the BW case. ∎
Finally, we can put all the previous lemmas together and prove the main theorem.
Theorem 1 (restated from the paper).
Any simple circuit on $N$ sites with periodic boundary conditions is equivalent to exactly one of the $N-1$ canonical simple circuits having Floquet propagator $F_{q,r}$, where $(q,r)\in Q_{N}$.
Proof.
The fact that $F_{q,r}$ are simple circuits follows from Lemma S3 and the number of them follows from Lemma S4.
According to Lemma S7, $C(F_{q,r})$ is different for all allowed $(q,r)$ and a given $N$. Since according to Lemma S1 $C$ is conserved under equivalence transformations and, according to Eq. (29), is different for all $F_{p}$, different $F_{q,r}$ are equivalent to different $F_{p}$ (according to Theorem S1, they must be equivalent to some $F_{p}$). Moreover, since there are $(N-1)$ $F_{q,r}$ and $(N-1)$ $F_{p}$, every $F_{p}$ must be equivalent to some $F_{q,r}$, i. e. there exists a bijective equivalence mapping between $F_{q,r}$ and $F_{p}$.
Given any simple circuit $F$, according to Theorem S1, we know $F\cong F_{p}$ for some $p$. We have shown that $F_{p}\cong F_{q,r}$ for some $q,r$. Therefore, $F\cong F_{q,r}$ for some $q,r$. ∎
In contrast to the proof of Theorem S1, the proof of Theorem 1 is not constructive, i. e. it cannot be used as an algorithm to transform a given $F$ to the equivalent $F_{q,r}$. The most convenient way to do this in practice is via the invariant $C(F)$ defined in Eq. (28), which can be numerically calculated in linear time in $N$. One can then calculate $C(F)$ and $C(F_{q,r})$ for the all $(N-1)$ allowed $(q,r)$ in quadratic time. Finally, $F$ must be equivalent to the $F_{q,r}$ with the same value of $C$. As an example, the values of $C(F_{q,r})$ are included in Fig. S1.